40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jun 26 16:53:38 2020, gautam, Update, Electronics, RFPD characterization WFS1.pdfWFS2.pdfbuildupMons.pdf
    Reply  Mon Jun 29 15:56:02 2020, gautam, Update, Electronics, RFPD characterization 
    Reply  Sun Jul 5 18:25:42 2020, rana, Update, Electronics, WFS characterization WFS_attenOff.pdf
       Reply  Sun Jul 12 22:40:35 2020, gautam, Update, Electronics, WFS characterization - old SURF report 
       Reply  Sat Aug 29 20:07:48 2020, rana, Update, Electronics, WFS characterization 
          Reply  Sat Aug 29 22:10:09 2020, gautam, Update, Electronics, WFS characterization 
          Reply  Tue Sep 15 17:04:43 2020, gautam, Update, Electronics, DC adaptors delivered 
Message ID: 15439     Entry time: Mon Jun 29 15:56:02 2020     In reply to: 15433
Author: gautam 
Type: Update 
Category: Electronics 
Subject: RFPD characterization 

A more comprehensive report has been uploaded here. I'll zip the data files and add them there too. In summary:

  1. There are several problems with the WFS heads
    • Some attenuators don't seem to work. This could be a problem with the Acromag BIO, or with the relay on the head itself.
    • The measured transimpedance at 29.5 MHz is much lower than expected. We expect ~50 kohms with no attenuation, and 5 kohms without. I measure 100 ohm - 2 kohm with the attenuation disabled, and ~200 ohms with it enabled.
    • Quadrant #3 on both WFS heads behaves differently from the others. There is also evidence of a 200 MHz oscillation for quadrant 3.
    • For some reason, there is a relative minus sign between the TFs measured for the WFS and for the RFPDs. I don't understand where this is coming from - all the OpAmps in the LSC PDs and WFS heads are configured as non-inverting, so why should there be a minus sign? Is this indicative of the polarity of the LEMO output being somehow flipped?
  2. POX 11 photodiode does not have a notch at 22 MHz.
  3. AS55 resonance appears to have shifted closer to 60 MHz, would benefit from a retuning. But the notches appear fine.
  4. PDA10CF photodiode used as the POP22/POP110 readback appears broken in some strange way. As shown in the linked document, a spare PDA10CF in the lab has a much more reasonable response, so I am going to switch out the POP22/POP110 diode with this spare.

I'll upload the data and analysis notebook + liso fit files to the wiki as well shortly. The data, a Jupyter notebook making the plots, and the LISO fit files have been uploaded here.

I didn't do it this time but it'd be nice to also do the noise measurement and get an estimate for the shot-noise intercept current.

Quote:

While I have the data, I will fit this and post a more complete report on the wiki.

ELOG V3.1.3-